# New Construction of Locally Perfect Nonlinear Functions with Application to Sequences Sets with Low Ambiguity Zone

### v2

@@ IV New LAZ sequence sets
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. Then
(i)  From Construction 1 (i), we have a LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$ is the number of sequences, $D=NK$ is the length of sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K$, $Z_{x}=min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, $\theta_{\text{max}}=K$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1$.
(ii)  From Construction 1 (ii), we have a LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$ is the number of sequences, $D=NK$ is the length of sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{x}=N$, $Z_{y}=p$, $p$ is the smallest prime factor of $K$, $\theta_{\text{max}}=K$, and $\hat{\theta}_{\text{max}}=K+N-1$.
@@
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. Then, from Construction 2, we have a LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$ is the number of sequences, $D=NK_{1}$ is the length of sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K_{1}-K+1$, $K\leq K_{1}$, $Z_{x}=min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, $\theta_{\text{max}}=K_{1}$, and $\hat{\theta}_{\text{max}}=K_{1}+Z_{x}-1$.
For the $(N,NK,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$-LAZ sequence set constructed in [1], the value of the parameter $N$ is only odd. But for the construction in Theorem 3, the value of parameter $N$, as a factor of $\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}$, can be chosen to be even. Thus the LAZ sequence set in Theorem 5 has more general and flexible parameters, and then one series of them is asymptotically optimal as shown in following corollary.
@@
Let $K=p^{e}$, where $p$ is an odd prime and $e\geq 1$, and let $N=\varphi(p^{e})=p^{e-1}(p-1)$. There exist LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N=\varphi(p^{e})$, $D=NK=p^{e}\varphi(p^{e})$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$, $\theta_{\text{max}}=K=p^{e}$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p^{e}+p-2$. Moreover, the series of LAZ sequence set is asymptotically optimal in periodic case when $p\rightarrow\infty$, and asymptotically optimal in aperiodic case when $e\geq 2$ and $p\rightarrow\infty$.
@@
The parameters $M$, $D$, $Z_{x}$, $Z_{y}$, $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$ can be seen directly from Theorem 3 and 5 by taking $s=R=1$. By Theorem 1, the YZLFLT bounds $\Delta$ and $\hat{\Delta}$ can be derived as follows:
$\left\{\begin{array}[]{ll}\Delta=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}=N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}},\\ \hat{\Delta}=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}=N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}},\end{array}\right.$
thus, we have
$\frac{\Delta}{\theta_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}}}{K}$
$=$ $\frac{N\sqrt{\frac{p-2}{N(p-1)-1}}}{\sqrt{K}}.$
Observe that if $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\Delta}{\theta_{\text{max}}}=\lim\limits_{p\to\infty}\sqrt{\frac{N}{K}}=1$ since $\frac{K}{N}=\frac{p^{e}}{p^{e-1}(p-1)}=\frac{p}{p-1}\rightarrow 1$. Furthermore, we have
$\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}}}{K+p-2}$
$=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-2)-(p-2)}{NKN(p-1)+N(p-2)(p-1)-(p+NK-2)}}}{K+p-2}.$
Observe that if $e\geq 2$ and $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}=\lim\limits_{p\to\infty}\frac{N\sqrt{K}}{K}\sqrt{\frac{NK(p-2)}{NKN(p-1)}}=\lim\limits_{p\to\infty}\sqrt{\frac{N}{K}}=1$ since $\hat{\theta}_{\text{max}}=p^{e}+p-2\rightarrow K$ and $\frac{K}{N}\rightarrow 1$. The proof of Corollary 1 is completed. ∎
Now we derive the case of twin-prime.
@@
Let $K=p(p+2)$, where $p$ and $p+2$ are odd prime, and let $N=\textrm{lcm}(p-1,p+1)=\frac{p^{2}-1}{2}$. There exist LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N=\frac{p^{2}-1}{2}$, $D=NK=\frac{p(p+2)(p^{2}-1)}{2}$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$, $\theta_{\text{max}}=K=p(p+2)$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p(p+2)+p-2$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$ and $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.

### v3

@@ IV Proposed LAZ sequence sets
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. Then:
(i)  A periodic LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,K)$ from Theorem 1 (i), where $M=N$ is the number of sequences, $D=NK$ is the length of the sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{x}=\min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, and $Z_{y}=K$.
(ii)  A periodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK,\Pi,K)$ from Theorem 1 (ii), where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{x}=N$, $Z_{y}=p$, and $p$ is the smallest prime factor of $K$.
Proof.
The results follow directly by applying the LPNFs given in Theorem 1 to Lemma 2, with the corresponding parameters $(M,D,\Pi,K)$ and $(N,NK,\Pi,K)$ specified in Theorem 3 (i) and (ii). ∎
Next, we present a specific LAZ sequence set in Theorem 3.
Corollary 1.
Let $K=p^{e}$, where $p$ is an odd prime, $e\geq 1$, and $N=\varphi(p^{e})=p^{e-1}(p-1)$. Let $\mathcal{S}$ denote a periodic LAZ sequence set from Theorem 3 above.
(i)  The $\mathcal{S}$ is an $(N,NK,\Pi,K)$ with $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$.
(ii)  The $\mathcal{S}$ is an $(N,NK,\Pi,K)$ with $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=p$, $Z_{x}=N=p^{e-1}(p-1)$.
Next, we discuss the optimality of the periodic LAZ sequence set presented in Corollary 1.
@@
Let $\mathcal{S}$ denote a periodic LAZ sequence set from Corollary 1 above. Then
(i)  the $\mathcal{S}$ presented in Corollary 1 (i) is asymptotically optimal with regard to the YZFLLT bound in Lemma 1 (i);
(ii)  the $\mathcal{S}$ presented in Corollary 1 (ii) is asymptotically optimal with regard to the YZFLLT bound in Lemma 1 (i).
Proof.
The parameters $M$, $D$, $Z_{x}$, $Z_{y}$, $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$ can be seen directly from Theorem 2 and 3 by taking $s=R=1$. By Theorem 1, the YZFLLT bound $\Delta$ can be derived as follows:
$\Delta$ $=$ $\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}$ (7)
$=$ $\sqrt{\frac{D^{2}MZ_{x}Z_{y}-D^{3}}{DZ_{y}(MZ_{x}-1)}}$
$=$ $\sqrt{\frac{D(MZ_{x}Z_{y}-D)}{Z_{y}MZ_{x}-Z_{y}}}$
$=$ $\sqrt{\frac{D(1-\frac{D}{MZ_{x}Z_{y}})}{1-\frac{1}{MZ_{x}Z_{y}}}}.$
(i)  Recall $M=N$, $D=NK$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$, then (7) yields
$\Delta$ $=$ $\sqrt{\frac{D(1-\frac{1}{p-1})}{1-\frac{1}{p^{e}p^{e-1}(p-1)^{2}}}}.$ (8)
Consider that if $p\rightarrow\infty$, then $\frac{K}{N}=\frac{p^{e}}{p^{e-1}(p-1)}=\frac{1}{1-\frac{1}{p}}\rightarrow 1$. Together with (8) and $\theta_{\text{max}}=K$, hence we infer
$\lim\limits_{p\to\infty}\frac{\theta_{\text{max}}}{\Delta}=\lim\limits_{p\to\infty}\sqrt{\frac{K^{2}}{D}}=\lim\limits_{p\to\infty}\sqrt{\frac{K}{N}}=1.$
(ii)  Recall $M=N$, $D=NK$, $Z_{y}=p$, $Z_{x}=N=p^{e-1}(p-1)$, it is easily checked that $\Delta$ is actually equal to (8), and so
$\lim\limits_{p\to\infty}\frac{\theta_{\text{max}}}{\Delta}=\lim\limits_{p\to\infty}\sqrt{\frac{K^{2}}{D}}=\lim\limits_{p\to\infty}\sqrt{\frac{K}{N}}=1.$
The proof of Theorem 4 is completed. ∎
Similarly, we will present some other common scenarios of $K$ in the Remark, i.e. $K$ is the product of twin primes or distinct odd primes.
Remark 3.
Case A: When $K=p(p+2)$, where $p$ and $p+2$ are odd prime, and suppose $N=\textrm{lcm}(p-1,p+1)=\frac{p^{2}-1}{2}$.
(i)  There exist a periodic LAZ sequence sets with parameters $(N,NK,\Pi,K)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$.
(ii)  There exist a periodic LAZ sequence sets with parameters $(N,NK,\Pi,K)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=p$, $Z_{x}=N=p^{e-1}(p-1)$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$.
Case B: When $K=p_{1}p_{2}\cdots p_{s}$, where $p_{1}<p_{2}<\cdots<p_{s}$ are odd primes, and suppose $N=E=\textrm{lcm}\{\varphi(p_{i})=p_{i}-1:1\leq i\leq s\}=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)}{2}$.
(i)  There exist a periodic LAZ sequence set with parameters $(N,NK,\Pi,K)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p_{1}p_{2}\cdots p_{s}$, $Z_{x}=p_{1}-1$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$ and $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
(ii)  There exist a periodic LAZ sequence set with parameters $(N,NK,\Pi,K)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=p_{1}$, $Z_{x}=N$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$.
The following Theorem presents the aperiodic LAZ set obtained via Theorem 1 and Lemma 2.
Theorem 5.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. We obtain:
(i)  An aperiodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK,\Pi,K+Z_{x}-1)$ from Theorem 1 (i), where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{x}=\min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, $Z_{y}=K$.
(ii)  An aperiodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK,\Pi,K+Z_{x}-1)$ from Theorem 1 (ii), where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{x}=N$, $Z_{y}=p$, $p$ is the smallest prime factor of $K$.
Proof.
The results follow directly by applying the LPNFs given in Theorem 1 to Lemma 2, with the corresponding parameters $(M,D,\Pi,K+Z_{x}-1)$ and $(N,NK,\Pi,K+Z_{x}-1)$ specified in Theorem 5 (i) and (ii). ∎
@@
Let $K=p^{e}$, where $p$ is an odd prime, $e\geq 1$, and $N=\varphi(p^{e})=p^{e-1}(p-1)$. Let $\mathcal{S}$ denote an aperiod LAZ sequence set from Theorem 5 above. Then
(i)   the $\mathcal{S}$ is an $(N,NK,\Pi,K+p-2)$ with $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$.
(ii)  the $\mathcal{S}$ is an $(N,NK,\Pi,K+N-1)$ with $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=p$, $Z_{x}=N=p^{e-1}(p-1)$.
The next Theorem analyzes the optimality of the aperiodic case.
Theorem 6.
Let $\mathcal{S}$ denote an aperiodic LAZ sequence set from Corollary 2 above. Then
(i)  the $\mathcal{S}$ presented in Corollary 2 (i) is asymptotically optimal under the YZFLLT bound in Lemma 1 (ii) if $e\geq 2$;
(ii)  the optimality factor of the $\mathcal{S}$ presented in Corollary 2 (ii) is $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=2$ under the YZFLLT bound in Lemma 1 (ii), where $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
@@
The parameters $M$, $D$, $Z_{x}$, $Z_{y}$, $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$ can be seen directly from Theorem 2 and 3 by taking $s=R=1$. By Theorem 1, the YZFLLT bound $\Delta$ can be derived as follows:
$\hat{\Delta}$ $=$ $\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}$ (9)
$=$ $\sqrt{\frac{D^{2}(MZ_{x}Z_{y}-D-Z_{x}+1)}{Z_{y}(MZ_{x}-1)(D+Z_{x}-1)}}$
$=$ $\sqrt{\frac{D(MZ_{x}Z_{y}-D-Z_{x}+1)}{Z_{y}(MZ_{x}-1)(1+\frac{Z_{x}}{D}-\frac{1}{D})}}$
$=$ $\sqrt{\frac{D(1-\frac{D}{MZ_{x}Z_{y}}-\frac{1}{MZ_{y}}+\frac{1}{MZ_{x}Z_{y}})}{(1-\frac{1}{MZ_{x}})(1+\frac{Z_{x}}{D}-\frac{1}{D})}}.$
(i)  Recall $M=N$, $D=NK$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$, then (9) yields
$\hat{\Delta}$ $=$ $\sqrt{\frac{D(1-\frac{1}{p-1}-\frac{1}{p^{e}p^{e-1}(p-1)}+\frac{1}{p^{e}p^{e-1}(p-1)^{2}})}{(1-\frac{1}{p^{e-1}(p-1)^{2}})(1+\frac{1}{p^{e}p^{e-1}}-\frac{1}{p^{e}p^{e-1}(p-1)})}}.$ (10)
we observe that if $e\geq 2$ and $p\rightarrow\infty$, then (10) implies
$\lim\limits_{p\to\infty}\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}=\lim\limits_{p\to\infty}\sqrt{\frac{(K+p-2)^{2}}{D}}=\lim\limits_{p\to\infty}\sqrt{\frac{(K+p-2)^{2}}{NK}}=1,$
since $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p^{e}+p-2\rightarrow p^{e}=K$ and $\frac{K}{N}=\frac{p^{e}}{p^{e-1}(p-1)}\rightarrow 1$.
(ii)  Recall $M=N$, $D=NK$, $Z_{y}=p$, $Z_{x}=N=p^{e-1}(p-1)$, (9) produces
$\hat{\Delta}$ $=$ $\sqrt{\frac{D(1-\frac{1}{p-1}-\frac{1}{p^{e}(p-1)}+\frac{1}{p^{e}p^{e-1}(p-1)^{2}})}{(1-\frac{1}{p^{2e-2}(p-1)^{2}})(1+\frac{1}{p^{e}}-\frac{1}{p^{e}p^{e-1}(p-1)})}}.$ (11)
But since $\hat{\theta}_{\text{max}}=K+Z_{x}-1=K+N-1$, combining (11) and $\frac{K}{N}=\frac{p^{e}}{p^{e-1}(p-1)}\rightarrow 1$, we obtain
$\lim\limits_{p\to\infty}\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$ $=$ $\sqrt{\frac{K^{2}+(N-1)^{2}+2K(N-1)}{NK}}$ (12)
$=$ $\sqrt{\frac{K^{2}}{NK}+\frac{N^{2}}{NK}-\frac{2N}{NK}+\frac{1}{NK}+\frac{2KN}{NK}-\frac{2K}{NK}}$
$=$ $\sqrt{\frac{K}{N}+\frac{N}{K}+2}$
$=$ $2$
The proof of Theorem 6 is completed. ∎
Remark 4.
Case A: When $K=p(p+2)$, where $p$ and $p+2$ are odd prime, and suppose $N=\textrm{lcm}(p-1,p+1)=\frac{p^{2}-1}{2}$. There exist an aperiodic LAZ sequence sets with parameters $(N,NK,\Pi,K)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$. Moreover, $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
Case B: When $K=p_{1}p_{2}\cdots p_{s}$, where $p_{1}<p_{2}<\cdots<p_{s}$ are odd primes, and suppose $N=E=\textrm{lcm}\{\varphi(p_{i})=p_{i}-1:1\leq i\leq s\}=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)}{2}$. There exist an aperiodic LAZ sequence set with parameters $(N,NK,\Pi,K+p_{1}-2)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p_{1}p_{2}\cdots p_{s}$, $Z_{x}=p_{1}-1$, $\theta_{\text{max}}=K=p_{1}p_{2}\cdots p_{s}$. Moreover, $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
In the above, we discuss the periodic and aperiodic LAZ sequence set by new LPNFs in Theorems 1 and their optimality. The following Theorem, we will state a series of periodic and aperiodic LAZ sequence set from Theorem 2.
Theorem 7.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. Then, from Theorem 2, we have:
(i)  A periodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK,\Pi,K_{1})$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K_{1}-K+2$, $K\leq K_{1}$, $Z_{x}=\min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$.
(ii)  An aperiodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK,\Pi,K_{1}+Z_{x}-1)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K_{1}-K+2$, $K\leq K_{1}$, $Z_{x}=\min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$.
Proof.
The results follow directly by applying the LPNFs given in Theorem 2 to Lemma 2, with the corresponding parameters $(N,NK,\Pi,K_{1})$ and $(N,NK,\Pi,K_{1}+Z_{x}-1)$ specified in Theorem 7 (i) and (ii). ∎
From the Throrem 7, the following result is immediate.
@@
Let $K=p^{e}$, where $p$ is an odd prime, $e\geq 1$, and let $N=\varphi(p^{e})=p^{e-1}(p-1)$. Then we obtain:
(i)  A periodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK_{1},\Pi,K_{1})$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K_{1}-K+2$, $Z_{x}=p-1$.
(ii)  An aperiodic LAZ sequence set $\mathcal{S}$ with parameters $(N,NK_{1},\Pi,K_{1}+p-2)$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K_{1}-K+2$, $Z_{x}=p-1$.
Theorem 8.
Let $\mathcal{S}$ denote a LAZ sequence set from Corollary 3 above. And let $\lim\limits_{p\to\infty}\sqrt{\frac{K_{1}}{N}}=1$ and $\lim\limits_{p\to\infty}\frac{K_{1}}{(p-1)(K_{1}-K+2)}=0$. Then
(i)  the $\mathcal{S}$ attains asymptotically optimal with respect to the periodic YZFLLT bound in Lemma 1 (i);
(ii)  the $\mathcal{S}$ attains asymptotically optimal with respect to the aperiodic YZFLLT bound in Lemma 1 (ii) if $e\geq 2$.